Urea is stronger base than other amides ?

Grade:12

1 Answers

Amit Sinha
9 years ago
The reason for this is this: Amides are weak bases due to resonance of the amide group: R–C(=O)–NH2 ? R–C(–O^-)=NH2+ In acetamide (R = CH3), the lone pair of nitrogen is not available for bonding with H^+. However, in urea (R = NH2), the lone pair on the second NH2 group is available some of the time for bonding with H^+, so it will be a stronger base than acetamide..